# The Self-Duality Equations on a Riemann Surface and Four-Dimensional Chern–Simons Theory

### v1

@@ 1.1 General Themes
is the usual Chern–Simons three-form. The theory (1.1) depends on the choice of the holomorphic volume form $\omega$ on $C$ which in general must have poles and possibly also zeroes. These zeroes and poles can be interpreted as defects and must be supplemented by boundary conditions. Once these are specified, they give a route to a Lagrangian formulation that incorporates the Lax pair of an integrable system into its Lagrangian formulation with the integrability of the Lax pair being the primary field equation arising from the four-dimensional Chern–Simons theory. Initial applications were to harmonic maps or sigma models, and Wess–Zumino–Witten [^6] These theories are sometimes also referred to as Wess–Zumino–Novikov–Witten models. (WZW) models [CostelloYamazaki201908]. It has gone on to play an important role in unifying a large class of both discrete [CostelloYamazakiWitten201809, CostelloYamazakiWitten201802] and continuous [CostelloYamazaki201908] integrable models in $1+1$ dimensions, with many more applications and developments such as [IshtiaqueMoosavianZhou201809, CostelloYagi201810, BittlestonSkinner201903, Vicedo201908, DelducLacroixMagroVicedo201909, Schmidtt201912, BeniniSchenkelVicedo202008, FukushimaSakamotoYoshida202003, CostelloStefanski202005, Tian202005, TianHeChen202007, Stedman202009, BittlestonSkinner202011, FukushimaSakamotoYoshida202105, Schmidtt202109, IshtiaqueMoosavianRaghavendranYagi202110, FukushimaSakamotoYoshida202112, BoujakhroutSaidi202204, BoujakhroutSaidi202207, LiniadoVicedo202301, BoujakhroutSaidiLaamaraDrissi202303, BoujakhroutSaidiLaamaraDrissi202304, BoujakhroutSaidiLaamaraDrissi202309, Khan202209, HeTianChen202105, Levine202309, Schmidtt202307, AshwinkumarSakamotoYamazaki202309, MoosavianYamazakiZhou202502, LacroixLevineWallberg202505, Schmidtt202508, SakamotoTateoYamazaki202509].
@@
Motivated by all the above, holomorphic Chern–Simons theories were introduced in the full 6d twistor space [Costello202004, BittlestonSkinner202011]. These give a Lagrangian underpinning to the original Ward construction [Ward:1977ta] adapted to give Lagrangian descriptions of the self-dual Yang–Mills equations in 4d in a gauge adapted to its realization as a 4d WZW model [LosevMooreNekrasovShatashvili199509]. Although holomorphic Chern–Simons formulations of the self-dual sector of maximally supersymmetric Yang–Mills have been around for some time [Witten:2003nn, Boels:2006ir] (or holomorphic BF theories in the non-supersymmetric case [Mason:2005zm]) these incorporate the antiselfdual (ASD) field also, rather than treating self-dual Yang–Mills as a self-contained theory. Because twistor space, $\mathbb{CP}^{3}$, is not Calabi–Yau, it has no non-singular holomorphic $3$-form so the theories of [Costello202004, BittlestonSkinner202011] used a holomorphic 3-form with poles. Like 4d Chern–Simons theory, for consistency these theories need boundary conditions at the poles. The corresponding symmetry reduction of these 6d Chern–Simons theories by the two real symmetries then directly yields the 4d Chern–Simons treatment of the reductions. The problem of fixed points and consequent singularities in the reduced twistor spaces mentioned above only arises if one attempts to reduce the twistor spaces as complex manifolds using the imaginary parts of the holomorphic symmetries. Thus 4d Chern–Simons theory is only partially complex but nevertheless captures the essentials of non-singular reductions of Ward’s twistor construction for these examples in a Dolbeault framework.
@@
A key motivation for this paper is to complete the realization of integrable models via four-dimensional Chern–Simons theory to the full range of 2d integrable systems obtained as reductions on a nondegenerate 2-plane given in [MasonWoodhouse199605, §6.2]. A particularly important example is that of Hitchin’s self-duality equations on a Riemann surface [Hitchin198707], whose realization in four-dimensional Chern–Simons theory is the main subject of the present work. They now play multiple fundamental roles across mathematics and physics [BeauvilleNarasimhanRamanan198907, BeilinsonDrinfeld1991, BeilinsonDrinfeld1996, HauselThaddeus200205, Ngo200406, Ngo200801, DonagiPantev200604, BershadskyJohansenSadovVafa199501, KapustinWitten200604, FrenkelWitten200710, DonagiWitten199510, GaiottoMooreNeitzke200907, NekrasovWitten201002, Gaiotto200911, NekrasovPestun201211]. Hitchin’s equations were first realized as the symmetry reduction of self-dual Yang–Mills equations in four dimensions of Euclidean signature, with two translation symmetries [Hitchin198707] and non-degenerate quotient. Remarkably, on reduction to two dimensions, the equations acquire two-dimensional conformal invariance and, as such, depend only on the complex structure of the surface $\Sigma$ on which they are defined. These equations are given by
@@
Another motivation in this work is to provide a Lagrangian understanding of canonical structures on the moduli spaces of solutions. In particular, Lagrangians lead to symplectic structures and Hamiltonians on the space of solutions of the corresponding equations, i.e., their moduli space. A particularly important case is when $\Sigma$ is taken to be a genus-$g\geq 2$ compact Riemann surface without boundary.Hitchin shows that the moduli space $\mathscr{M}_{\text{H}}(\Sigma,G)$ of solutions to (1.3) on $\Sigma$ has a hyperkähler structure, so that it has a metric with three distinguished metric-compatible, covariantly-constant complex structures denoted by $\mathcal{I},\mathcal{J},$ and $\mathcal{K}$ that satisfy the quaternion relations [Hitchin198707]. Lowering the index with the metric, gives three corresponding symplectic forms denoted by $\Omega_{\mathcal{I}},\Omega_{\mathcal{J}},$ and $\Omega_{\mathcal{K}}$, given in (2.51) and (2.58). These complex structures and symplectic forms naturally combine into a $\mathbb{CP}^{1}$-family of complex structures $\mathcal{J}_{\zeta}$ and corresponding symplectic forms parametrized by $\zeta\in\mathbb{CP}^{1}$ given explicitly by
$\Omega_{\mathcal{J}_{\zeta}}:=\mathbf{a}(\zeta,\bar{\zeta})\,\Omega_{\mathcal{I}}+\mathbf{b}(\zeta,\bar{\zeta})\,\Omega_{\mathcal{J}}+\mathbf{c}(\zeta,\bar{\zeta})\,\Omega_{\mathcal{K}},$ (1.4)
with
$\left(\mathbf{a}(\zeta,\bar{\zeta}),\mathbf{b}(\zeta,\bar{\zeta}),\mathbf{c}(\zeta,\bar{\zeta})\right):=\frac{1}{1+|\zeta|^{2}}\left(1-|\zeta|^{2},\mathrm{i}(\zeta-\bar{\zeta}),(\zeta+\bar{\zeta})\right)\,.$ (1.5)
@@
According to the choice of $\zeta$, $\mathscr{M}_{\text{H}}(\Sigma,G)$ has two different interpretations. In the first, in the complex structures $\pm\,\mathcal{I}$ corresponding to $\zeta=0,\infty$, it is the moduli space of Higgs bundles [^7] A Higgs bundle is a pair $(E,\varphi)$ where $E$ is a holomorphic vector bundle $E\to\Sigma$, and $\varphi\in\Omega^{1,0}_{\Sigma}(\text{End}(E))$ is an $\text{End}(E)$-valued holomorphic $1$-form, the $(1,0)$ part of $\phi$. satisfying a suitable stability condition. Assuming certain (semi-)stability conditions, $\mathscr{M}_{\text{H}}(\Sigma,G)\simeq T^{*}\mathscr{M}_{\text{Hol}}(\Sigma,G)$, where $\mathscr{M}_{\text{Hol}}$ denotes the moduli space of (semi-)stable holomorphic bundles on $\Sigma$, and $\simeq$ should be understood as a birational equivalence of the two spaces. Hitchin then shows that $\Omega_{\mathcal{J}}+i\Omega_{\mathcal{K}}$ defines a holmorphic symplectic form for the complex structure $\mathcal{I}$, and finds a complete set of $\mathcal{I}$-holomorphic commuting Hamiltonians realizing $\mathscr{M}_{\text{H}}(\Sigma,G)$ as the phase space of an algebraically-complete Hamiltonian integrable system [Hitchin198701], see §2.4 for more details. The second interpretation follows for all complex structures $\mathcal{J}_{\zeta}$ for $\zeta\neq 0,\infty$ where $\mathscr{M}_{\text{H}}(\Sigma,G)$ can be identified with the moduli space of complexified flat connections, also known as the character variety. The relationship between these two interpretations of $\mathscr{M}_{\text{H}}(\Sigma,G)$ is the subject of non-Abelian Hodge correspondence developed in [Donaldson198707, Corlette198801, Hitchin199207, Simpson198810, Simpson199212]. These constructions, including the definition of Higgs bundles, can be extended to noncompact surfaces with punctures equipped with parabolic structures [Simpson199007, Konno199304, Yokogawa199301, BodenYokogawa, Sabbah199905, BiquardBoalch200401, Boalch199506] (and also higher-dimensional Kähler manifolds [Simpson199212]). The actions of 4d Chern–Simons theory, which lead to $\Omega_{\mathcal{I}}$ and $\Omega_{\mathcal{J}_{\zeta}}$ will be presented in §2 and §3, respectively.
@@ 1.2 Contributions of the Paper
2d Action for the Hitchin’s Equations (§2.1).
@@
Here, $h:\Sigma\to G\setminus G_{\mathbb{C}}$, $\psi:\Sigma\to\text{Lie}(G)$, where $|\cdot|_{h}$ is defined in (2.8) and $S_{\mathrm{WZW}}[h]$ is the Wess–Zumino–Witten action, defined in (2.9). For $G=\text{SU}(n)$, $h\in\mathrm{Herm}_{0}$, the space of positive definite Hermitian matrices of unit determinant since we can write $h=f_{0}^{*}f_{0}$ for $f_{0}:\Sigma\to\mathrm{SL}(n,\mathbb{C})$ identifying the coset $\mathrm{SU}(n)\setminus\mathrm{SL}(n,\mathbb{C})$ with $h\in\mathrm{Herm}_{0}$. Using $\mathrm{d}_{\Sigma}=\partial+\bar{\partial}$ and the identification
@@
we show that the field equations and gauge freedoms of (1.6) coincide with the Hitchin’s equations (1.3) in a holomorphic gauge.
2d Action for the Hitchin’s Equations from 4d CS Theory (§2.32.4).
@@
Since $\omega$ has poles at $z=0,\infty$ we impose the boundary conditions
@@
Then if the field equations are imposed, there exists a gauge in which the new gauge field $\mathcal{A}$ can be written as
@@
so that $\omega(\zeta,\bar{\zeta})$ has zeroes at $z=0,\infty$ where we allow the following simple poles
@@
We furthermore impose the reality conditions (1.11). Then, we can reduce to a corresponding family of 2d actions $S_{\text{H},\zeta}$ parametrized by $\mathbb{CP}^{1}$ given in (3.10), which are gauge equivalent to (1.1) for various choices of $\zeta\in\mathbb{CP}^{1}$. The field equations of this action are then equivalent to the Hitchin’s equations. Finally, the symplectic structure on the space of solutions of the field equations of $S_{\text{H},\zeta}$ arising from this action formulation is $\Omega_{\mathcal{J}_{\zeta}}$, given in (1.4). ∎
@@
We highlight that there are (at least) two known actions that lead to Hitchin’s equations: one gives the Lax form using a three-dimensional topological-holomorphic version of BF theory [CaudrelierHarlandSinghVicedo202509, Thms. 4.1 and 4.3] (see also [VicedoWinstone202201] for a related work), and the second action uses a certain topological-holomorphic BF-type theory in four dimensions to derive Hitchin’s equations [Jarov202509, §7.1]. Such BF theories require the introduction of Lagrange multiplier fields and so do not give rise to canonical symplectic structures directly on the original solution space without Lagrange multipliers. Since the formulations of this paper are based on 4d Chern–Simons rather than BF theories, we obtain natural symplectic structures involving potentials for the Hitchin connection and Higgs field. This has been the main motivation for the present work. Our formulation also makes the role of twistor parameter $\zeta$ quite explicit, whereas the role of the twistor parameter is not so clear in [CaudrelierHarlandSinghVicedo202509, Jarov202509]. ∎
@@ 2 From 4d Chern–Simons Theory to Hitchin’s Equations
This section first presents an action for Hitchin’s equations in §2.1 and goes on to define a corresponding four-dimensional Chern–Simons theory as in (1.1) in §2.2. Then in §2.3 we give a proof of Proposition 1.1 that explain’s how the 2d action can be reduced from that in 4d. We show explicitly that in a local patch of $\Sigma$, the action (1.1) subject to appropriate boundary conditions at the poles of $\omega$ descends to our 2d Lagrangian that gives rise to the Hitchin equations. We go on to compute the symplectic structure on the space of solutions of the field equations in §2.4, and show that it coincides with the canonical symplectic structure on $\mathscr{M}_{\text{H}}(\Sigma,G)$ associated with the complex structure $\mathcal{I}$. We go on to explain how to construct the Hitchin Hamiltonians on the moduli space from four-dimensional Chern–Simons theory.
2.1 2d Action for the Hitchin’s Equations
@@
where $F_{A}=\mathrm{d}A+A\wedge A$ is the curvature of $D$ and $\widetilde{\varphi}\in\Omega^{0,1}_{\Sigma}(\mathrm{End}(E))$. More generally, we can complexify the system, so that the connection $A$ is complex and $\varphi^{*}$ replaced by an indpendent $\widetilde{\varphi}$, with reality subsequently imposed by the condition that there exists a unitary gauge in which $A=-A^{*}$ and $\widetilde{\varphi}=\varphi^{*}$, the Hermitian conjugate to $\varphi$. These equations form a 2d integrable system with the Lax pair defined by the connection
@@
In order to write a second-order action of standard type, we introduce local potentials for our fields $(A,\varphi,\varphi^{*})$. The connection defines a holomorphic structure on $E$ as there is no integrability condition on a Riemann surface. Then we can find a local holomorphic frame but this will not in general be unitary, and the Hermitian metric will take the general form
@@
In this latter Wess–Zumino term, a factor of $\mathrm{i}$ is included for $h$ Hermitian, and as usual, an extension of $h$ from $\Sigma\times 0$ to $\Sigma\times[0,1]$ has been taken that is fixed to be the identity at $\Sigma\times 1$. The third homotopy group of the space of unit determinant positive definite Hermitian matrices is trivial, ($\text{SL}(n,\mathbb{C})/\mathrm{SU}(n)$ is contractible) so there is no ambiguity in the second term, and its normalization does not need to be fixed. [^9] This is essentially the two-dimensional version of the Donaldson action described in [Donaldson198501], see [MasonWoodhouse199605, §3.4] for an alternative local formula. If we were to change the reality conditions to those for a harmonic map, then we would need to impose the usual normalization that leads to an ambiguity of integral multiples of $2\pi\mathrm{i}$ in this term.
It can be seen directly that the action yields the Hitchin equations. The $\widetilde{\psi}$ variation of the second term of (2.8) yields $\bar{\partial}\varphi=0$, as defined in (2.7) which is equivalent to $\bar{\partial}_{A}\varphi=0$ in this holomorphic gauge. On the other hand, the $h$-variation of $S_{\mathrm{WZW}}$ yields the curvature $F_{A}$ and the variation of the second term of (2.8) yields the commutator of the Higgs field with its Hermitian conjugate as required. Furthermore, the action is invariant under the choice of $\psi$ as the second term is the squared norm of the Higgs field. It is also invariant under the choice of a holomorphic frame, as the variation of $h$ just leads to gauge covariant equations of motion.
@@
We now explain the version of 4d Chern–Simons theory for the Hitchin equations. Consider the action (1.1) on $\Sigma\times\mathbb{CP}^{1}$, i.e., $C=\mathbb{CP}^{1}$; we will for the most part work locally on a Riemann surface $\Sigma$. The gauge field $\mathsf{A}$ of four-dimensional Chern–Simons theory defines a connection $D_{\mathsf{A}}$ on a bundle $E\rightarrow\Sigma\times\mathbb{CP}^{1}$ with structure group $G=\mathrm{SL}(n,\mathbb{C})$. We will assume that $E$ is topologically trivial on the $\mathbb{CP}^{1}$ factors. Although $\Sigma$ and $\mathbb{CP}^{1}$ denote the topological and holomorphic planes of the theory, our boundary conditions will also require a complex structure on $\Sigma$ so we introduce complex coordinates $(w,z)$, with $w\in\mathbb{C}$ a local coordinate on $\Sigma$ and $z$ on $\mathbb{CP}^{1}$. Since $\mathsf{A}$ is wedged against $\omega$, which is a multiple of $\mathrm{d}z$, a meromorphic $(1,0)$-form on $\mathbb{CP}^{1}$, $\mathsf{A}$ is a partial connection lying $\Omega^{1}_{\Sigma}\oplus\Omega^{0,1}_{\mathbb{CP}^{1}}$. The 1-form $\omega$ must have poles and can also have zeroes and this data will be supplemented with appropriate boundary conditions on $\mathsf{A}$ at the zeros and poles of $\omega$.

### v2

@@ 1.1 General Themes
is the usual Chern–Simons three-form. The theory (1.1) depends on the choice of the holomorphic volume form $\omega$ on $C$ which in general must have poles and possibly also zeroes. These zeroes and poles can be interpreted as defects and must be supplemented by boundary conditions. Once these are specified, they give a route to a Lagrangian formulation that incorporates the Lax pair of an integrable system into its Lagrangian formulation, with the integrability of the Lax pair being the primary field equation arising from the four-dimensional Chern–Simons theory. Initial applications were to harmonic maps or sigma models, and Wess–Zumino–Witten [^6] These theories are sometimes also referred to as Wess–Zumino–Novikov–Witten models. (WZW) models [CostelloYamazaki201908]. It has gone on to play an important role in unifying a large class of both discrete [CostelloYamazakiWitten201809, CostelloYamazakiWitten201802] and continuous [CostelloYamazaki201908] integrable models in $1+1$ dimensions, with many more applications and developments such as [IshtiaqueMoosavianZhou201809, CostelloYagi201810, BittlestonSkinner201903, Vicedo201908, DelducLacroixMagroVicedo201909, AshwinkumarTan201910, Schmidtt201912, FukushimaSakamotoYoshida202003, CostelloStefanski202005, Tian202005, TianHeChen202007, BeniniSchenkelVicedo202008, Stedman202009, BittlestonSkinner202011, CaudrelierStoppatoVicedo202012, Zucchini202101, FukushimaSakamotoYoshida202105, Schmidtt202109, IshtiaqueMoosavianRaghavendranYagi202110, FukushimaSakamotoYoshida202112, BoujakhroutSaidi202204, BoujakhroutSaidi202207, IshtiaqueZhou202211, LiniadoVicedo202301, BoujakhroutSaidiLaamaraDrissi202303, Khan202209, HeTianChen202105, Levine202309, Schmidtt202307, AshwinkumarSakamotoYamazaki202309, MoosavianYamazakiZhou202502, LacroixLevineWallberg202505, Schmidtt202508, SakamotoTateoYamazaki202509].
@@
Motivated by all the above, holomorphic Chern–Simons theories were introduced in the full 6d twistor space [Costello202004, BittlestonSkinner202011]. These give a Lagrangian underpinning to the original Ward construction [Ward:1977ta] adapted to give Lagrangian descriptions of the self-dual Yang–Mills equations in 4d in a gauge adapted to its realization as a 4d WZW model [LosevMooreNekrasovShatashvili199509]. Although holomorphic Chern–Simons formulations of the self-dual sector of maximally supersymmetric Yang–Mills have been around for some time [Witten:2003nn, Boels:2006ir] (or holomorphic BF theories in the non-supersymmetric case [Mason:2005zm]) these incorporate the antiselfdual (ASD) field also, rather than treating self-dual Yang–Mills as a self-contained theory. Because twistor space, $\mathbb{CP}^{3}$, is not Calabi–Yau, it has no non-singular holomorphic $3$-form so the theories of [Costello202004, BittlestonSkinner202011] used a holomorphic 3-form with poles. Like 4d Chern–Simons theory, for consistency, these theories need boundary conditions at the poles. The corresponding symmetry reduction of these 6d Chern–Simons theories by the two real symmetries then directly yields the 4d Chern–Simons treatment of the reductions. The problem of fixed points and consequent singularities in the reduced twistor spaces mentioned above only arises if one attempts to reduce the twistor spaces as complex manifolds using the imaginary parts of the holomorphic symmetries. Thus, 4d Chern–Simons theory is only partially complex but nevertheless captures the essentials of non-singular reductions of Ward’s twistor construction for these examples in a Dolbeault framework.
@@
A key motivation for this paper is to complete the realization of integrable models via four-dimensional Chern–Simons theory to the full range of 2d integrable systems obtained as reductions on a nondegenerate 2-plane given in [MasonWoodhouse199605, §6.2]. A particularly important example is that of Hitchin’s self-duality equations on a Riemann surface [Hitchin198707], whose realization in four-dimensional Chern–Simons theory is the main subject of the present work. They now play multiple fundamental roles across mathematics and physics [BeauvilleNarasimhanRamanan198907, BeilinsonDrinfeld1991, BeilinsonDrinfeld1996, HauselThaddeus200205, Ngo200406, Ngo200801, DonagiPantev200604, BershadskyJohansenSadovVafa199501, KapustinWitten200604, FrenkelWitten200710, DonagiWitten199510, GaiottoMooreNeitzke200907, NekrasovWitten201002, Gaiotto200911, NekrasovPestun201211]. Hitchin’s equations were first realized as the symmetry reduction of self-dual Yang–Mills equations in four dimensions of Euclidean signature, with two translation symmetries [Hitchin198707] and a non-degenerate quotient. Remarkably, on reduction to two dimensions, the equations acquire two-dimensional conformal invariance and, as such, depend only on the complex structure of the surface $\Sigma$ on which they are defined. These equations are given by
@@
Another motivation in this work is to provide a Lagrangian understanding of canonical structures on the moduli spaces of solutions. In particular, Lagrangians lead to symplectic structures and Hamiltonians on the space of solutions of the corresponding equations, i.e., their moduli space. A particularly important case is when $\Sigma$ is taken to be a genus-$g\geq 2$ compact Riemann surface without boundary. Hitchin shows that the moduli space $\mathscr{M}_{\text{H}}(\Sigma,G)$ of solutions to (1.3) on $\Sigma$ has a hyperkähler structure, so that it has a metric with three distinguished metric-compatible, covariantly-constant complex structures denoted by $\mathcal{I},\mathcal{J},$ and $\mathcal{K}$ that satisfy the quaternion relations [Hitchin198707]. Lowering the index with the metric gives three corresponding symplectic forms denoted by $\Omega_{\mathcal{I}},\Omega_{\mathcal{J}},$ and $\Omega_{\mathcal{K}}$, given in (2.51) and (2.58). These complex structures and symplectic forms naturally combine into a $\mathbb{CP}^{1}$-family of complex structures $\mathcal{J}_{\zeta}$ and corresponding symplectic forms parametrized by $\zeta\in\mathbb{CP}^{1}$ given explicitly by
$\Omega_{\mathcal{J}_{\zeta}}:=\mathbf{a}\,\Omega_{\mathcal{I}}+\mathbf{b}\,\Omega_{\mathcal{J}}+\mathbf{c}\,\Omega_{\mathcal{K}},$ (1.4)
where
$\left(\mathbf{a},\mathbf{b},\mathbf{c}\right):=\frac{1}{1+|\zeta|^{2}}\left(1-|\zeta|^{2},\mathrm{i}(\zeta-\bar{\zeta}),(\zeta+\bar{\zeta})\right)\,.$ (1.5)
@@
According to the choice of $\zeta$, $\mathscr{M}_{\text{H}}(\Sigma,G)$ has two different interpretations. In the complex structures $\pm\,\mathcal{I}$ corresponding to $\zeta=0,\infty$, it is the moduli space of Higgs bundles [^7] A Higgs bundle is a pair $(E,\varphi)$ where $E$ is a holomorphic vector bundle $E\to\Sigma$, and $\varphi\in\Omega^{1,0}_{\Sigma}(\text{End}(E))$ is an $\text{End}(E)$-valued holomorphic $1$-form, the $(1,0)$ part of $\phi$. satisfying a suitable stability condition. Assuming certain (semi-)stability conditions, $\mathscr{M}_{\text{H}}(\Sigma,G)\simeq T^{*}\mathscr{M}_{\text{Hol}}(\Sigma,G)$, where $\mathscr{M}_{\text{Hol}}$ denotes the moduli space of (semi-)stable holomorphic bundles on $\Sigma$, and $\simeq$ should be understood as a birational equivalence of the two spaces. Hitchin then shows that $\Omega_{\mathcal{J}}+i\Omega_{\mathcal{K}}$ defines a holomorphic symplectic form for the complex structure $\mathcal{I}$, and finds a complete set of $\mathcal{I}$-holomorphic commuting Hamiltonians realizing $\mathscr{M}_{\text{H}}(\Sigma,G)$ as the phase space of an algebraically completely Hamiltonian integrable system [Hitchin198701], see §2.4 for more details. The second interpretation follows for all complex structures $\mathcal{J}_{\zeta}$ for $\zeta\neq 0,\infty$ where $\mathscr{M}_{\text{H}}(\Sigma,G)$ can be identified with the moduli space of complexified flat connections, also known as the character variety. The relationship between these two interpretations of $\mathscr{M}_{\text{H}}(\Sigma,G)$ is the subject of non-Abelian Hodge correspondence developed in [Donaldson198707, Corlette198801, Hitchin199207, Simpson198810, Simpson199212]. These constructions, including the definition of Higgs bundles, can be extended to noncompact surfaces with punctures equipped with parabolic structures [Simpson199007, Konno199304, Yokogawa199301, BodenYokogawa, Sabbah199905, BiquardBoalch200401, Boalch199506] (and also higher-dimensional Kähler manifolds [Simpson199212]). The actions of 4d Chern–Simons theory, which lead to $\Omega_{\mathcal{I}}$ and $\Omega_{\mathcal{J}_{\zeta}}$ will be presented in §2 and §3, respectively.
@@ 1.2 Contributions of the Paper
2d Action for Hitchin’s equations (§2.1).
@@
Here, $h:\Sigma\to\text{Herm}_{0}$, $\psi:\Sigma\to\text{Lie}(G)$, where $|\cdot|_{h}$ is defined in (2.8) and $S_{\mathrm{WZW}}[h]$ is the Wess–Zumino–Witten action, defined in (2.9). For $G=\text{SU}(n)$, $h\in\mathrm{Herm}_{0}$, the space of positive definite Hermitian matrices of unit determinant, since we can write $h=f_{0}^{*}f_{0}$ for $f_{0}:\Sigma\to\mathrm{SL}(n,\mathbb{C})$ identifying the coset $\mathrm{SU}(n)\setminus\mathrm{SL}(n,\mathbb{C})$ with $h\in\mathrm{Herm}_{0}$. Using $\mathrm{d}_{\Sigma}=\partial+\bar{\partial}$ and the identification
@@
we show that the field equations and gauge freedoms of (1.6) coincide with Hitchin’s equations (1.3) in a holomorphic gauge.
2d Action for Hitchin’s equations from 4d CS Theory (§2.32.4).
@@
Since $\omega$ has poles at $z=0,\infty$, we impose the boundary conditions
@@
Then, if the field equations are imposed, there exists a gauge in which the new gauge field $\mathcal{A}$ can be written as
@@
so that $\omega(\zeta,\bar{\zeta})$ has zeroes at $z=0,\infty$ where we allow the following boundary condition
@@
We furthermore impose the reality conditions (1.11). Then, we can reduce to a corresponding family of 2d actions $S_{\text{H},\zeta}$, given in (3.10), parametrized by $\mathbb{CP}^{1}$, which are gauge equivalent to (1.1) for various choices of $\zeta\in\mathbb{CP}^{1}$. The field equations of this action are then equivalent to Hitchin’s equations. Finally, the symplectic structure on the space of solutions of the field equations of $S_{\text{H},\zeta}$ arising from the four-dimensional Chern–Simons theory is $\Omega_{\mathcal{J}_{\zeta}}$, given in (1.4). ∎
@@
We highlight that there are (at least) two known actions relating to Hitchin’s equations: One is a Lagrange multiplier type action obtained by compactifying 4d mixed topological-holomorphic BF theory to 2d dimensions [Jarov202509, §7.1]. This is the 2d counterpart of the Chalmers–Siegel [Chalmers:1996rq] action for self-dual Yang–Mills theory, whereas the functional (1.6) is the 2d counterpart of the 4d WZW model, or Donaldson functional. We also note that the gauged 4d Chern–Simons set-ups of [ColeCullinanHoareLiniadoThompson202407, §4] yield 2d integrable field theories coupled to a Hitchin subsector. On the other hand, a variational formulation of the Hitchin integrable system based on a three-dimensional topological-holomorphic version of BF theory has been constructed in [CaudrelierHarlandSinghVicedo202509, Thms. 4.1 and 4.3] (see also [VicedoWinstone202201, Cole:2025zmq] for related works). We emphasise that the Hitchin integrable system lives on the moduli space of solutions to Hitchin’s equations, so although both functionals have a natural description in terms of mixed topological-holomorphic field theory, they do not appear to be immediately related. ∎
@@ 2 From 4d Chern–Simons Theory to Hitchin’s Equations
This section first presents an action for Hitchin’s equations in §2.1 and goes on to define a corresponding four-dimensional Chern–Simons theory as in (1.1) in §2.2. Then in §2.3 we give a proof of Proposition 1.1 that explains how the 2d action can be reduced from that in 4d. We show explicitly that in a local patch of $\Sigma$, the action (1.1) subject to appropriate boundary conditions at the poles of $\omega$ descends to our 2d Lagrangian that gives rise to the Hitchin equations. We go on to compute the symplectic structure on the space of solutions of the field equations in §2.4, and show that it coincides with the canonical symplectic structure on $\mathscr{M}_{\text{H}}(\Sigma,G)$ associated with the complex structure $\mathcal{I}$. We go on to explain how to construct the Hitchin Hamiltonians on the moduli space from four-dimensional Chern–Simons theory.
2.1 2d Action for Hitchin’s equations
@@
where $F_{A}=\mathrm{d}A+A\wedge A$ is the curvature of $D$ and $\widetilde{\varphi}\in\Omega^{0,1}_{\Sigma}(\mathrm{End}(E))$. More generally, we can complexify the system, so that the connection $A$ is complex and $\varphi^{*}$ replaced by an independent $\widetilde{\varphi}$, with reality subsequently imposed by the condition that there exists a unitary gauge in which $A=-A^{*}$ and $\widetilde{\varphi}=\varphi^{*}$, the Hermitian conjugate to $\varphi$. These equations form a 2d integrable system with the Lax pair defined by the connection
@@
In order to write a second-order action of standard type, we introduce local potentials for our fields $(A,\varphi,\varphi^{*})$. The connection defines a holomorphic structure on $E$ as there is no integrability condition on a Riemann surface. Then we can find a local holomorphic frame, but this will not, in general, be unitary, and the Hermitian metric will take the general form
@@
In this latter Wess–Zumino term, a factor of $\mathrm{i}$ is included for $h$ Hermitian, and as usual, an extension of $h$ from $\Sigma\times 0$ to $\Sigma\times[0,1]$ has been taken that is fixed to be the identity at $\Sigma\times 1$. The third homotopy group of the space of unit determinant positive definite Hermitian matrices is trivial, ($\text{SL}(n,\mathbb{C})/\mathrm{SU}(n)$ is contractible) so there is no ambiguity in the second term, and its normalization does not need to be fixed. [^9] If we were to change the reality conditions to those for a harmonic map, then we would need to impose the usual normalization that leads to an ambiguity of integral multiples of $2\pi\mathrm{i}$ in this term. This is essentially the two-dimensional version of the action in [Donaldson198501] for Hermitian Yang-Mills due to Donaldson, who already observes its equivalence with the a WZW-type action in [Donaldson:1992vc], see also [Simpson198810] for applications to Higgs bundles and also [MasonWoodhouse199605, §3.4] for an alternative local formula.
It can be seen directly that the full action yields the Hitchin equations. The $\widetilde{\psi}$ variation of the second term of (2.8) yields $\bar{\partial}\varphi=0$, as defined in (2.7), which is equivalent to $\bar{\partial}_{A}\varphi=0$ in this holomorphic gauge. On the other hand, the $h$-variation of $S_{\mathrm{WZW}}$ yields the curvature $F_{A}$ and the variation of the second term of (2.8) yields the commutator of the Higgs field with its Hermitian conjugate as required. Furthermore, the action is invariant under the choice of $\psi$ as the second term is the squared norm of the Higgs field. It is also invariant under the choice of a holomorphic frame, as the variation of $h$ just leads to gauge covariant equations of motion.
@@
We now explain the version of 4d Chern–Simons theory for the Hitchin equations. Consider the action (1.1) on $\Sigma\times\mathbb{CP}^{1}$, i.e., $C=\mathbb{CP}^{1}$; we will, for the most part, work locally on a Riemann surface $\Sigma$. The gauge field $\mathsf{A}$ of four-dimensional Chern–Simons theory defines a connection $D_{\mathsf{A}}$ on a bundle $E\rightarrow\Sigma\times\mathbb{CP}^{1}$ with structure group $G=\mathrm{SL}(n,\mathbb{C})$. We will assume that $E$ is topologically trivial on the $\mathbb{CP}^{1}$ factors. Although $\Sigma$ and $\mathbb{CP}^{1}$ denote the topological and holomorphic planes of the theory, our boundary conditions will also require a complex structure on $\Sigma$, so we introduce complex coordinates $(w,z)$, with $w\in\mathbb{C}$ a local coordinate on $\Sigma$ and $z$ on $\mathbb{CP}^{1}$. Since $\mathsf{A}$ is wedged against $\omega$, which is a multiple of $\mathrm{d}z$, a meromorphic $(1,0)$-form on $\mathbb{CP}^{1}$, $\mathsf{A}$ is a partial connection lying $\Omega^{1}_{\Sigma}\oplus\Omega^{0,1}_{\mathbb{CP}^{1}}$. The 1-form $\omega$ must have poles and can also have zeroes, and this data will be supplemented with appropriate boundary conditions on $\mathsf{A}$ at the zeros and poles of $\omega$.